OpenStudy (anonymous):

the length of a rectangle is 3m less than twice the width, and the area of the rectangle is 14m^2. find the dimensions of the rectangle.

OpenStudy (anonymous):

Let the width be 'w' Length = 2w - 3 Area of the rectangle = 14 (2w - 3)(w) = 14 2w^2 - 3w - 14 = 0 2w^2 + 4w - 7w - 14 = 0 2w(w + 2) - 7(w + 2) = 0 (2w - 7)(w + 2) = 0 w = 7/2, -2 Since width cannot be negative, w = 7/2 length = 4

OpenStudy (anonymous):

thank you, so much! I could not figure out what I did wrong, but I see the problem now.